OpenStudy (user123):

can we simplify 3(x+1)^2+4(x+1)^3 further

OpenStudy (user123):

can this be 12(x+1)^5

OpenStudy (3mar):

May I help?

OpenStudy (user123):

Yes please

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Your expression is \[3(x+1)^2+4(x+1)^2\] Is that right?

OpenStudy (user123):

yes.

OpenStudy (3mar):

ok \[3x+4x=??\] \[7xy+8xy=??\] \[2x^2+6x^2=???\]

OpenStudy (user123):

7x 15 xy 8x^2

OpenStudy (3mar):

Great!

OpenStudy (3mar):

Take (x+1) as a common factor, what you get?

OpenStudy (user123):

7(x+1)^5

OpenStudy (user123):

?

OpenStudy (3mar):

I mean \[(x+1)^2[(3+4(x+1)]=(x+1)^2[3+4x+4]=(x+1)^2[4x+7]\] Is it clear now?

OpenStudy (user123):

kind of. i dont know why you did it that way.

OpenStudy (3mar):

in order to simplify the parentheses and make it simpler

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (user123):

ah okay.

OpenStudy (3mar):

Any more questions?

OpenStudy (user123):

thank yoi so much for the help.

OpenStudy (3mar):

Thank you for learning!! Any Help... Anytime.